Tstaff will be upgrading the OS on the backend servers providing database service, for a number of our production services, starting at 07:30 on Tuesday, March 15th. This will affect: the CS website, CS email lists, our problem ticket system (RT), CS wikis, instant messaging service, and a variety of Tstaff utlitilies. Assuming all ...
09:57: We experienced some authentication issues with our VPN server. The process has been restarted and we are investigatng the root cause.
17:03: It looks like there is a libgcrypt bug that crops up when you use openVPN and authenticate against LDAP over TLS, which is exactly our setup. This bug results in ...
***The workstation reboot is scheduled for tomorrow morning, February 18th, at 6am.***
Note: Grid nodes and virtual machines will be rebooted after coordinating ...
Google posted to their online security blog earlier today about an interesting discovery one of their engineers made while trying to debug an SSH problem. After some investigation, they determined it was actually due to a glibc blug that could be exploited. The patches rolled out in coordination with this blog post. Unfortunately, it does ...
Early this morning we migrated from using physical hardware to virtualized machines to handle internal mail delivery. We expect there to be no impact on our users, but are letting everyone know in case you notice some change in behavior. These servers are responsible for all delivery of email to problem, our issue tracking system ...
We added another disk tray with 14.55TB of disks to the GPFS Data filesystem. We are running a disk restripe to rebalance the Data filesystem. This could potentially affect filesystem performance.
Tstaff will be performing a department-wide reboot of Linux desktop machines at 5:30am tomorrow, Wednesday, January 20th, to apply an urgent security patch. Please be sure to log out of your machines before then.
A major vulnerability was announced today in versions 3.8 and higher of the Linux kernel. The details may be ...
The department experienced a GPFS blip shortly after 10am this morning. The outage become pregressively worse and lasted about 10 minutes in total.
For those interested in more details, one of the quorum and manager nodes for the cluster experienced a system hard drive failure a few days ago. After replacing the hard drive and ...
We rolled out some changes over lunch today that affect how our departmental workstations reach the rest of the Brown campus and the internet. While we don't expect there to be any issues, we wanted to let everyone know in case they experienced connectivity issues.
For those interested in the network details, read on ...
- January 2016
- February 2016
- March 2016
- April 2016
- May 2016
- June 2016
- July 2016
- August 2016
- September 2016
- October 2016
- November 2016
- December 2016